OpenStudy (rainbow_rocks03):

What is the measure of ∠ABC? A. 76° B. 86° C. 44° D. 104°

OpenStudy (rainbow_rocks03):



OpenStudy (rainbow_rocks03):

I think the answer is 104.D.

OpenStudy (rainbow_rocks03):

@rishavraj am I correct?

Mehek (mehek14):

no ur wrong

OpenStudy (anonymous):

no.

OpenStudy (rainbow_rocks03):

oh

Mehek (mehek14):

all the angles add up to 180 56 + 48 = 104 180 - 104 = answer

OpenStudy (anonymous):

add 46 and 58 and then subract by 180

Mehek (mehek14):

46 + 58 my bad

OpenStudy (rainbow_rocks03):

so I was correct right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

r u in k12?

OpenStudy (rainbow_rocks03):

oopss hang on

rishavraj (rishavraj):

sum of all angles of a triangle = 180 so 58 + 46 + angle ABC = 180 angle ABC = 180 - 104

Mehek (mehek14):

no u weren't you added 48 + 56 = 104

Mehek (mehek14):

56 + 48*

OpenStudy (rainbow_rocks03):

yah I added 48+56 to get 104

OpenStudy (anonymous):

once u get 104 subtract it by 180

Mehek (mehek14):

thats the sum of 2 angles you want to find the third angle

Mehek (mehek14):

the third angle has to be 104 + x = 180

OpenStudy (rainbow_rocks03):

104-180=-76

OpenStudy (rainbow_rocks03):

that's wrong

rishavraj (rishavraj):

why u getting negative ??? o_O

Mehek (mehek14):

180 - 104

OpenStudy (anonymous):

180-104

OpenStudy (anonymous):

a

OpenStudy (anonymous):

ru in k12 @rainbow_rocks03

OpenStudy (rainbow_rocks03):

I got negative because @thebro55 said to subtract 104-180

OpenStudy (rainbow_rocks03):

I said yes

OpenStudy (anonymous):

i said 104 then subtract it from 180

OpenStudy (anonymous):

and what grade? i am in 8th. i had this math class a while back